OpenStudy (anonymous):

If the simple interest on a sum of 2 years at 10%p.a is Rs.1000 and the same amount is borrowed for 2 years at 8%p.a and compounded annually, then find the compound interest and theamout to be paid after 2 years.

OpenStudy (anonymous):

a typing mistake = theamout should be the amount

OpenStudy (anonymous):

on a sum of 2 years????

OpenStudy (anonymous):

P * T * R S.I. = -------- 100 P * 2 * 10 1000 = --------- 100 1000 * 100 ---------- = P 20 so P = 5000 Now put this P in the formula for compound interest and calculate. If u need help, let me know...

OpenStudy (anonymous):

do i need to find the amount after the principle

OpenStudy (anonymous):

Use this formula for calculating amount [ R ]^n A = P [1 + ---- ] [ 100 ] where n = no. of years then C.I = A - P

OpenStudy (anonymous):

i did the answer in this way, is it correct P * T * R S.I. = -------- 100 P * 2 * 10 1000 = --------- 100 1000 * 100 ---------- = P 20 P = 5000 A = 1000 + 5000 = 6000 According to the sum ow A becomes P P = Rs.6000 [ R ]^n A = P [1 + ---- ] [ 100 ] THEN A came as Rs.6,998.4 AND CI = 6,998.4 - 6000 = Rs.19948.4 Is it correct

OpenStudy (anonymous):

"ow" moeans "now"

OpenStudy (anonymous):

no this is wrong first we find the P using the S.I. formula i.e. P = Rs. 5000 the question says "the same amount" which means the same P is borrowed on C.I. so we have to use the same P i.e. Rs. 5000 in the formula for C.I.

OpenStudy (anonymous):

THANKS FOR THE HELP

OpenStudy (anonymous):

u r welcome........☺